OpenStudy (anonymous):

Given the quadratic function f(x)=x^2-6x+9, find a value of x such that f(x)=25. Show all work, please.

OpenStudy (saifoo.khan):

x^2-6x+9 = 25 solve.

OpenStudy (saifoo.khan):

x^2 -6x -16

OpenStudy (saifoo.khan):

Solve thsi!

OpenStudy (saifoo.khan):

this*

OpenStudy (anonymous):

what goes into -6 that goes into -16. I have nothing other than 1 and -16

OpenStudy (anonymous):

I have (x + )(x- )

OpenStudy (anonymous):

f(x) = 25 means x² - 6x + 9 = 25 x² - 6x + 9 - 25 = 0 x² - 6x - 16 = 0 x² -8x + 2x - 16 = 0 x(x - 8) + 2(x - 8) = 0 (x + 2)(x - 8) = 0 this is possible if either x+2 = 0 or x-8 = 0 if x + 2 = 0 ==> x = -2 if x - 8 = 0 ==> x = 8 so f(x) = 25 for x = -2 and x = 8

OpenStudy (saifoo.khan):

Hakirat, is 100% corrct.

OpenStudy (anonymous):

thanks saifoo :D